Lauren wants to mix 5 liters of 7% milk with skim- milk(0%fat) to produce a mixture of 2 46/47% milk. How much skim-milk should Lauren add?
Answer by stanbon(75887)

Lauren wants to mix 5 liters of 7% milk with skim- milk(0%fat) to produce a mixture of 2 46/47% milk. How much skim-milk should Lauren add?
---------
Comment: Follow the fat
Also 2 46/47% = 0.0297
----------
7% solution DATA:
Amt = 5 liters ; amount of fat = 0.07*5 = 0.35 liters
------------
skim solution DATA:
Amt = x liters ; amount of fat = 0*x = 0 liters
--------------
Mixture DATA:
Amt = 5+x liters ; amount of fat = 0.0297(5+x)=0.149+0.0297x liters
-------------
EQUATION:
fat + fat = fat
0.35 + 0 = 0.149+0.0297x
0.32x = 0.149
x = 0.465 liters (amount of skim that should be added)
==============
Cheers,
Stan H.